OpenStudy (anonymous):

CAN ANYONE HELP ME TO FIND THA INTEGRAL SIN(2(3.14)t/T-X) Dt

OpenStudy (anonymous):



OpenStudy (anonymous):

here. T-x in the denominator is a constant. so, it will just be the integral of sin 2 pi t which is 1/(2 pi) * cosine (2 pi* t) replace the limit values.

OpenStudy (anonymous):

\[\sin 2( \pi t) = 2\sin (\pi t) \cos (\pi t)\] This might help

OpenStudy (anonymous):